Two numbers differ by 3. Four times the lesser diminished by three times the greater is 7. Find the numbers.
Answer by checkley77(12569)

x-y=3
x=y+3
4y-3x=7
4y-3(y+3)=7
4y-3y-9=7
y=7+9
y=16 answer.
x-16=3
x=3+16
x=19 answer.
Proof:
4*16-3*19=7
64-57=7
7=7